The healthcare provider prescribes cefazolin (Ancef) 125 mg Q6H for a child weighing

50 lb. Safe dose is 25-50 mg/kg in 4 divided doses. Available is 125 mg/mL.

Is this dose safe? Please circle one as your answer:

safe, but ineffective

safe

not safe.

If safe, how many mL/dose would the nurse give? ___________mL

PLEASE PLEASE SHOW WORK IN DIMENSIONAL ANALYSIS ONLY THANK YOU
Answer #1

ANSWER.

* Safe dose of the child.

          Ordered dose = 125 mg Q6H or four times a day.

           Weight of the patient = 50 lb.

Convert lb to kg for calculation

            1 lb = 0.454 kg

     So

         Weight of the patient = 50 * 0.454

                                            = 22.7 kg.

    Safe dose = 25 - 50 mg/kg in 4 divided doses

    Safe dose for this child per day = 25 * 22.7 - 50 * 22.7

                                          = 567.5 mg - 1135 mg per day in 4 divided doses.

     Safe dose for this child per dose = 567.5/4 - 1135/4

                                                          = 141.8 mg to 283.7 mg per dose.

            Ordered dose = 125 mg per dose.

* So the ordered dose is SAFE,but ineffective for this child.